# WisdoM[Uncaptioned image]: Improving Multimodal Sentiment Analysis by Fusing Contextual World Knowledge

### 12 Jan 2024 v1

@@ 5 Experiment
ALMT Zhang et al. (2023)* 83.73 83.98 83.73
-w/ WisdoM 89.62 (+5.89) 89.51 (+5.53) 89.54 (+5.81)
@@ 5.1 Experimental Settings
Datasets
For aspect-level task, our two benchmark datasets are Twitter2015 and Twitter2017 Yu and Jiang (2019). Twitter2015 and Twitter2017 consist of multimodal tweets. Each multimodal tweet consists of text, an image posted alongside the tweet, aspects within the tweet, and the sentiment of each aspect. Each aspect is given a label from the set {negative, neutral, positive}. For sentence-level task, we evaluate our WisdoM on MSED Jia et al. (2022) dataset. MSED is a multimodal and multi-task dataset, which contains 9,190 text-image pairs.
Method Twitter2015 MSED
Acc. F1 Mac-F1
MMICL 76.0 72.7 86.2
 + context 69.5 72.1 86.8
 + Contextual Fusion 77.3 74.2 88.9
Table 3: Ablation study of context and its wise fusion module. We first only incorporate “context” and subsequently introduce the “contextual fusion” module.
Models
To demonstrate that our approach generalized across various architectures and sizes, we experimented on MMICL (14B) Zhao et al. (2023), LLaVA-v1.5 (13B) Liu et al. (2023a), mPLUG-Owl2 (8.2B) Ye et al. (2023), AoM (105M) Zhou et al. (2023), ALMT (112.5M) Zhang et al. (2023).
Evaluation Metrics
For aspect-level tasks, we use Accuracy (Acc) and F1 following previous studies Khan and Fu (2021); Ju et al. (2021); Ling et al. (2022); Zhou et al. (2023). For the sentence-level task, we adopt precision, recall, and macro-F1 (F1) as evaluation metrics.
Implement Details
For task-specific methods Zhou et al. (2023); Zhang et al. (2023), we employ the same inference procedures as their original works. For general-purpose methods Liu et al. (2023a); Zhao et al. (2023); Ye et al. (2023), we compute the likelihood that a LVLM generates the content of this choice given the question. We select the choice with the highest likelihood as model’s prediction. For fair comparison, these models are finetuned on target datasets. The learning rate is grid searched in {1e-5, 2e-5, 7.5e-5, 1e-4}, and batch size is in {16, 32, 256}. We optimize our model with AdamW Loshchilov and Hutter (2018).
5.2 Hyperparameter Selection
In stage 3, Contextual fusion has two major hyperparameters $\beta$ and $\alpha$. To systematically study the impact of $\beta$ and $\alpha$, we first fix $\alpha=0.3$ and search different configurations for the $\beta$. Then we fix $\beta$ to the optimal value in the Twitter2015 dev set and search for $\alpha$. As shown in Fig. 4, we find that $\beta$ values within the range of $[0.4,0.5]$ demonstrate strong performance for both models. We observe that as the $\beta$ increases, indicating the introduction of more context information, all models exhibit a substantial decrease. We hypothesize that this decline may be caused by introducing excessive noise within the context. Fig. 5 shows the trend on Twitter2015 dev set for different $\alpha$ values. We note that performance remains strong for both models when $\alpha$ values are in the range of $[0.3,0.4]$. In addition, we observe no notable decrease in trends as $\alpha$ increases. One possible explanation for this phenomenon is that the introduced context proves to be effective for samples with low confidence, but its impact diminishes for samples with high confidence. For a fairer comparison, we fixed $\alpha=0.3$, and $\beta$ is the optimal value on the dev set for testing.
@@ 5.3 Main Results
Evaluation for Aspect-level MSA Task
We compare against SOTA on Twitter2015 and 2017. Table 1 shows the performance of aspect-level tasks. Our WisdoM achieves consistent and significant improvement on four models across two datasets. The WisdoM brings max 2.9% and 2.3% F1-gains on Twitter2015 and Twitter2017 respectively. The results show that our method has a clear advantage.
Evaluation for Sentence-level MSA Task
As shown in Table 2, the best F1 on MSED, our method WisdoM with LLaVA-v1.5 is 90.48, outperforming LLaVA-v1.5 (88.75), mPLUG-Owl2 (87.98), MMICL (86.17) and ALMT (83.73). The significant improvement on the SOTA MSA method ALMT brings 5.81% F1-gains. In addition, we observe a more pronounced enhancement in the effect on small-scale models, whereas the impact is less prominent on larger models. This suggests that contextual world knowledge is particularly crucial for small models in MSA.
@@ 5.4 Ablation Study
Effect of Different Modules.
To better understand the role of each module in our method WisdoM, Table 3 presents the ablation results of the gradual addition of different components. Compared with the baseline (MMICL), adding context results in performance degradation. We found that context containing irrelevant information leads to degradation of performance on Twitter2015. The further addition of Contextual Fusion brings a significant improvement of 2.1 F1 score.
Effect of Different Fusion Strategies.
In Table. 5.4, we explore different fusion strategies experiments. We conducted experiments involving $mean(P_{i},\hat{P}_{i})$ and $max(P_{i},\hat{P}_{i})$. We observe that we can achieve a marginal improvement with these straightforward fusion strategies. Furthermore, we calculate the Jensen-Shannon (JS) divergence of $P_{i}$ with the uniform distribution to serve as the fusion weight, i.e., $\beta$. However, JS divergence is influenced by the class with the smallest probability, introducing an impractical level of difficulty.

### 20 Feb 2024 v2

@@ 5 Experiments
ALMT Zhang et al. (2023b)* 83.73 83.98 83.73
-w/ WisdoM 89.92 (+6.19) 90.14 (+6.16) 90.01 (+6.28)
@@ 5.1 Experimental Settings
Datasets.
For aspect-level tasks, our two benchmark datasets are Twitter2015 and Twitter2017 Yu and Jiang (2019). Twitter2015 and Twitter2017 are comprised of multimodal tweets, where each tweet incorporates textual content, an accompanying image, aspects contained within the tweet, and the sentiment associated with each aspect. Each aspect is assigned a label from the predefined set {negative, neutral, positive}. For sentence-level tasks, we evaluate our WisdoM on a multimodal and multi-task dataset MSED Jia et al. (2022), containing 9,190 text-image pairs.
Models.
To demonstrate WisdoM generalizes across architectures and sizes, we experimented on MMICL (14B) Zhao et al. (2023), LLaVA-v1.5 (13B) Liu et al. (2023a), mPLUG-Owl2 (8.2B) Ye et al. (2023), AoM (105M) Zhou et al. (2023), ALMT (112.5M) Zhang et al. (2023b).
Evaluation Metrics.
For aspect-level tasks, we use Accuracy (Acc.) and macro-F1 (F1) following previous studies Khan and Fu (2021); Ju et al. (2021); Ling et al. (2022); Zhou et al. (2023). For the sentence-level task, we adopt precision, recall, and macro-F1 (F1) as evaluation metrics.
@@ 5.2 Main Results
Results of Aspect-Level MSA Task.
We compare against advanced aspect-level MSA methods on Twitter2015 and Twitter2017, and report the results on Table 1. We show that our WisdoM achieves consistent and significant improvement on four models across two datasets. The WisdoM brings max 2.9% and 2.3% F1-gains on Twitter2015 and Twitter2017 respectively, showing that our method has a clear advantage.
Results of Sentence-Level MSA Task.
As shown in Table 2, notably, our WisdoM upon LLaVA-v1.5 achieves the new SOTA F1 score: 90.48%, outperforming LLaVA-v1.5 (88.75%), mPLUG-Owl2 (87.98%), MMICL (86.17%) and ALMT (83.73%), consistently. The most significant improvement is achieved on ALMT (112.5M), where we bring an encouragingly 6.28% F1 gain, suggesting that contextual world knowledge is particularly crucial for small models in MSA.
Method Twitter2017 MSED
Acc. F1 F1
MMICL
Baseline 74.1 74.1 86.2
[…]
 + CF 75.7 (+1.6) 75.7 (+1.6) 88.9 (+2.7)
LLaVA-v1.5
Baseline 74.6 74.3 88.8
[…]
 + CF 75.6(+1.0) 75.3 (+1.0) 90.5 (+1.7)
Table 3: Ablation study of context and its wise fusion module. “CF” denotes our Contextual Fusion. We first only incorporate “context” and subsequently introduce the “contextual fusion” module.
Figure 4: Comparative winning rates of Our Context v.s. RAG-based methods on Twitter2015 and Twitter2017 benchmarks. We can see that our contexts are better than the knowledge provided by RAG and PKG.
Method Twitter2015 Twitter2017
Acc. F1 Acc. F1
RAG 75.1 71.0 71.9 70.7
PKG 76.2 72.4 72.8 71.5
Our Context 76.3 72.7 73.7 73.5
Table 4: Comparative results of context generated by RAG, PKG, and our stage 2. We incorporate the contexts on LLaVA-v1.5 directly.
@@ 5.3 Ablation Study
To better understand the role of each module in our method WisdoM, Table 3 presents the ablation results of the gradual addition of different components. Compared with the baselines (MMICL and LLaVA-v1.5), only adding context results in a slight performance degradation (-0.23% average F1 score), while with the help of our proposed Context Fusion mechanism, we achieve a consistent and significant improvement (+1.75% average F1 score). Through (error) case studies in Appendix C.2, we found that containing irrelevant information in the original context leads to bad performance, showing the necessity of further context fusion mechanism.
Analyzing Effects of Context.
To further analyse the effect of context, we compare our context generated in stage 2 with the document retrieved by RAG Lewis et al. (2020) and knowledge generated by PKG Luo et al. (2023), collectively termed as “context” for simplicity. The assessment focuses on the context’s pertinence to a given image $v$ and sentence $s$, alongside its applicability in MSA tasks. We employ the LLM-based metric, i.e., LLM-as-a-Judge Chiang et al. (2023a) to quantify the quality of context. Specifically, we craft a prompt for GPT-4V OpenAI (2023) to compare our context with that provided by RAG and PKG. The detailed experimental settings can be found in Appendix A.7. As shown in Fig. 4, our context “Our Context” significantly beats the RAG and PKG counterparts, demonstrating its superiority. We provide the examples of context provided by different methods in Appendix C.1. Besides analyzing the contexts’ pertinence of different methods, we report their downstreaming performance on MSA tasks in Table 4. Clearly, the MSA performance with our context is the best. The results above illustrate that our method can provide more precise context, thus bringing better MSA performance.
Analyzing Effects of Contextual Fusion.
[…]
@@
5.4 When and Why Does Our Method Work?
To better understand when and why our method works? we conduct extensive analysis to provide the following insights:
Figure 5: Effects of different types of world knowledge. We analyse the effect of different types of world knowledge by applying WisdoM to AoM. The orange dash line and blue dash line represent the F1-score of vanilla AoM on Twitter2015&2017 respectively.
Figure 6: Comparison of context ($S_{c\rightarrow p}$) and input’s ($S_{i\rightarrow p}$) correlation to the final prediction across layers in LLaVA-v1.5 on Twitter2015. High score means a strong correlation with final decision-making.
Figure 7: Visualizing of error rate for hard samples ($\delta\leq 0.3$) on Twitter2015 benchmark.
World Knowledge Enhances MSA, while Domain-related Knowledge is more Helpful.
We explore the effects of different types of world knowledge from 11 perspectives (artistic, biographical, etc.), upon AoM. Fig. 5 shows that 1) nearly all types of extra knowledge enhance the MSA performance, 2) historical knowledge significantly enhances MSA on Twitter datasets. We conjecture that this is because tweets often convey sentiment through historical references, and domain-related world knowledge is more helpful. To verify our hypothesis, we conduct experiments on Financial PhraseBank (FPB, Malo et al., 2014) dataset and Twitter financial sentiment validation dataset Magic (2022), and find that financial and historical knowledge are the two types of world knowledge with the greatest gain (bringing 4.2% and 2.1% improvements on average F1 scores, respectively), confirming our conjunction. Details can be refer to Appendix B.5.
Context is Dominant for Prediction.
To draw a clearer picture of the information flow for MSA, we calculate $S_{c\rightarrow p}$ and $S_{i\rightarrow p}$ to represent the mean significance of information flow from context (c) and original input (i) to the prediction labels (p) respectively (Simonyan et al. (2014); Wang et al. (2023a), detailed in Appendix A.8). Fig. 6 reveals that the significance of the information flow from context to the prediction label is remarkably high than its input counterpart, suggesting that context outweighs image and sentence when making the final prediction. Results of other datasets show identical trends and can be found in Appendix B.6.
WisdoM Effectively Reduces the Uncertainty of Hard Samples.
To further explore how our WisdoM affects the hard samples, we visualize the error rate within high entropy in Fig. 7. After integrating WisdoM, the error rate is significantly decreased compared with the baseline (AoM and LLaVA-v1.5), demonstrating that our WisdoM effectively reduces the uncertainty of hard samples and improves performance.
6 Conclusion
In this paper, we propose a simple and effective plug-in framework WisdoM[Uncaptioned image] to enhance the ability of MSA. WisdoM[Uncaptioned image] follows three stages: Prompt Templates Generation, Context Generation, and Contextual Fusion. Firstly, we employ ChatGPT to generate prompt templates, enabling the large vision-language model to produce pertinent contextual world knowledge (referred to as context) derived from both the image and sentence. Subsequently, we incorporate this context using Contextual Fusion, minimizing the introduction of noise in the process. We empirically demonstrated the effectiveness and universality of the WisdoM on a series of widely used benchmarks.
Limitation
Our work has several potential limitations. Although our experiments revealed the enhanced performance of the multimodal sentiment analysis model with the introduction of context, the adaptive incorporation of context requires further exploration. Additionally, there is a need for further research into introducing models capable of handling additional modalities (e.g., speech).
Ethics Statement
We take ethical considerations very seriously. This paper focuses on improving multimodal sentiment analysis by making the most of large vision-language models. All experiments are conducted on open datasets and the findings and conclusions of this paper are reported accurately and objectively. Thus, we believe that this research will not pose ethical issues.
Appendix A Implementation Details
A.1 Details of Datasets
In this work, we conduct experiments covering aspect-level (Twitter2015 and Twitter2017) and sentence-level (MSED) MSA benchmarks. We present the statistics of all datasets in Table 6. Then, each dataset is described as:
Twitter2015 & 2017.
Twitter2015 and Twitter2017 datasets encompass multimodal tweets, wherein each tweet comprises textual content, an accompanying image, embedded aspects, and corresponding sentiment annotations for each aspect. These aspects are categorized into labels from the set {negative, neutral, positive}.
MSED.
MSED comprises 9190 pairs of text and images sourced from diverse social media platforms, including but not limited to Twitter, Getty Images, and Flickr.
Twitter2015 Twitter2017 MSED
#Train #Dev #Test #Train #Dev #Test #Train #Dev #Test
Negative 368 149 113 416 144 168 1939 308 613
Neutral 1883 679 607 1638 517 573 1664 294 569
Positive 928 303 317 1508 515 493 2524 419 860
Total 3179 1122 1037 3562 1176 1234 6127 1021 2042
Table 6: Dataset statistics.
A.2 Model Details
To verify the effectiveness of our WisdoM, we apply it within two standard frameworks for modeling the MSA tasks. For task-specific framework, we conduct experiment using the AoM (105M, Zhou et al., 2023) and ALMT (112.5M, Zhang et al., 2023b). For general-purpose framework, we experiment on mPLUG-Owl2 (8.2B, Ye et al., 2023), LLaVA-v1.5 (13B, Liu et al., 2023a) and MMICL (14B, Zhao et al., 2023). The detailed model information is listed in Table 7.
Model Model Type Source
AoM Task-specific https://github.com/SilyRab/AoM
ALMT Task-specific https://github.com/Haoyu-ha/ALMT
LLaVA-v1.5 General-purpose https://huggingface.co/liuhaotian/llava-v1.5-13b
mPLUG-Owl2 General-purpose https://huggingface.co/MAGAer13/mplug-owl2-llama2-7b
MMICL General-purpose https://huggingface.co/BleachNick/MMICL-Instructblip-T5-xxl
Table 7: Information of all models used in this study.
A.3 Training Details
For a fair comparison, we fine-tune models introduced in § 5.1 on target datasets. We optimize our model with AdamW Loshchilov and Hutter (2018). The learning rate is grid searched in {1e-5, 2e-5, 7.5e-5, 1e-4}, and batch size is in {16, 32, 256}. It should be noted that contexts are not incorporated during the training stage.
A.4 Inference Details
For task-specific methods Zhou et al. (2023); Zhang et al. (2023b), we employ the same inference procedures as their original works. For general-purpose methods Liu et al. (2023a); Zhao et al. (2023); Ye et al. (2023), we compute the likelihood that an LVLM generates the content of this choice given the question. We select the choice with the highest likelihood as the model’s prediction.
A.5 Prompt Templates of World Knowledge
In Table 8, we list the prompt templates which are provided by ChatGPT in prompt templates generation stage. In the context generation stage, we replace “[x]” with sentence $s_{i}$ and a special image token (e.g., “<image>”) is inserted at the beginning of the prompt template.
A.6 Pseudo Code for Contextual Fusion
Algorithm 1 provides the pseudo-code of Contextual Fusion. The simplicity of our method requires only a few lines of code.
# p_o: numpy array, represents $P_{i}$
# p_c: numpy array, represents $\hat{P}_{i}$
# alpha: float, the threshold of choosing hard sample
# beta: float, interpolation coefficient
def contextual_fusion(p_o, p_c, alpha, beta):
# step1 : calculate delta
delta = 2 * np.max(p_o) + np.min(p_o) - 1
if delta > alpha:
return p_o
# step2 : calculate the final prediction
p_f = p_o + beta * (p_c - p_o)
return p_f
Algorithm 1 Python-style pseudo-code for Contextual Fusion.
Type name
Prompt template
Artistic
Identify and discuss any artistic movements or styles that influenced the creation of the image. Explore how the artist’s choice of style aligns with or deviates from prevalent artistic trends of the time. Sentence: [x].
Biographcial
Delve into the backgrounds of individuals associated with the image and text. Explore the biographies of artists, authors, or other relevant figures, and discuss how their life experiences shaped the creation and interpretation of the work. Sentence: [x]
Character
Focus on characters within the image and sente. Analyze their personalities, relationships, and potential character development. Discuss how the visual and textual elements contribute to character portrayal. Sentence: [x]
Cultural
Explore how the image and sentence reflect or represent aspects of a particular culture. Discuss the cultural significance, traditions, or values implied by the elements in the image and sentence. Sentence: [x]
Environmental
Examine the environmental elements within the image and sentence, discussing ecological factors, environmental changes, or the relationship between human activities and the depicted setting. Sentence: [x]
Historical
Give you an image and sentence, you can provide historical context, important events, and relevant background information related to the image and sentence. Sentence: [x]
Literary
Conduct a literary analysis of the sentence, exploring themes, symbolism, and narrative techniques. Discuss how the words complement or contrast with the visual elements in the image. Sentence: [x]
Political
Examine the political during the time the image and text were created. Discuss any political events, movements, or ideologies that may have influenced the content and tone of the work. Sentence: [x]
Scientific
Investigate the scientific elements within the image, delving into discoveries, advancements, or breakthroughs related to the subject matter mentioned in the sentence. Sentence: [x]
Social
Investigate the image and text as a form of social commentary. Analyze how the work reflects or critiques social issues, norms, or inequalities prevalent at the time of creation. Sentence: [x]
Financial
Give you a sentence and image, you should provide related financial knowledge. Sentence: [x]
Table 8: Example of the prompt template generated in stage 1 and used in stage 2.
A.7 RAG Experiment Setup
In § 5.3, we compare our WisdoM with two RAG-based methods: (1) a naive RAG Lewis et al. (2020), which initially searches for relevant documents related to a given question and then employs a generator to predict an answer; (2) PKG Luo et al. (2023), an advanced RAG method that incorporates a knowledge-guided module, allowing for information retrieval without modifying the parameters of language models. The experimental setting is described as below.
Knowledge Sources.
The source of knowledge for our experiments is the Wikipedia-Image-Text (WIT) dataset Srinivasan et al. (2021), which has published in 2021. This dataset comprises images from Wikipedia, along with their alt-text captions and contextualized text passages.
Methods.
For naive RAG, we use off-the-shelf Contriever-MSMARCO Izacard et al. (2022) as the textual retriever and CLIP-ViT as the visual retriever. Specifically, we utilize CLIP-ViT and Contriever-MSMARCO for encoding query $q$ and knowledge source, and employ Maximum Inner Product Search (MIPS, Guo et al., 2020) to find the five nearest neighbors (knowledge) to $q$ within the knowledge source. For PKG, we use LLaVA-v1.5 (13B) as knowledge-guided module fine-tuning on WIT and then generate the knowledge according to the image $v$ and sentence $s$. Subsequently, we directly predict sentiment polarity by incorporating knowledge.
Evaluation Setting.
To evaluate the relevance of the context to a specific image $v$ and sentence $s$, we employ LLM-based metric, i.e., LLM-as-a-Judge Chiang et al. (2023a). Specifically, a prompt is crafted for GPT-4V OpenAI (2023) to assess the winning rates of our context in comparison to those derived from RAG-based methods. The detailed prompt can be found in Table 9. We also evaluate performance on MSA tasks.
Evaluation Prompt
**System**: In this task, you will be asked to compare the relevance of two paragraphs to determine which one is more pertinent to the provided source sentence and image and benefits the sentiment analysis task the most. There are three options for you to choose from:
1. Context1 is better. If you think Context 1 is more relevant to the source sentence and image and benefits the sentiment analysis task.
2. Context2 is better. If you think Context 2 is more relevant to the source sentence and image and benefits the sentiment analysis task.
3. Context1, Context2 are the same: If you think Context1, Context2 have the same relevance to the source sentence and image, then choose this option.
*Your answer is a JSON DICT that has one key: answer. For example: {"answer": "x. Context x is better."}**
*INPUT**
Source Sentence: “[s]”
Context1: “[x1]”
Context2: “[x2]”
*OUTPUT**
Table 9: The Evaluation Prompt we used for GPT-4V. [s] represents the input sentence. [x1] and [x2] represent the context generated by different methods.
A.8 Calculation of $S_{c\rightarrow p}$ and $S_{i\rightarrow p}$
To measure the significance of context and original input (i.e., image and sentence), we use $S_{c\rightarrow p}$ and $S_{i\rightarrow p}$ for highlighting critical token interactions. Following previous work Simonyan et al. (2014); Wang et al. (2023a), we use the Taylor expansion Michel et al. (2019) to calculate the score for each element of the attention matrix:
$I_{l}=\sum_{h}|A_{h,l}^{\mathsf{T}}\frac{\partial\mathcal{L}(x)}{\partial A_{h,l}}|.$ (3)
Here, $A_{h,l}$ represents the value of the attention matrix of the $h$-th attention head in the $l$-th layer, $x$ is the input, and $\mathcal{L}(x)$ is the loss function. We calculate the saliency matrix $I_{l}$ for the $l$-th layer by averaging across all attention heads. $I_{l}(k,j)$ denotes the importance of the information flow from the $j$-th word to the $k$-th word. We propose two quantitative metrics based on $I_{l}$. The definitions of the two quantitative metrics are below.
$S_{c\rightarrow p}$, the mean significance of information flow from context($c$) to the prediction label ($p$).
$S_{c\rightarrow p}=\frac{\sum_{(k,j)\in C_{cp}}I_{l}(k,j)}{|C_{cp}|},$ (4)
[…]
$S_{i\rightarrow p}$, the mean significance of information flow from image ($v$) and sentence ($t$) to the prediction label ($p$).
$S_{i\rightarrow p}=\frac{\sum_{(k,j)\in C_{ip}}I_{l}(k,j)}{|C_{ip}|},$ (5)
$C_{i,p}=\{(i,p):i\in[v,t]\}.$
$S_{c\rightarrow p}$ and $S_{i\rightarrow p}$ indicate the intensity of information aggregation onto the prediction label. A high $S$ demonstrates strong information for final decision-making.
Appendix B Additional Experimental Results
B.1 Hyperparameter Selection
In stage 3, Contextual fusion has two major hyperparameters interpolation coefficient $\beta$ and uncertain threshold $\alpha$. To systematically study the impact of $\beta$ and $\alpha$, we first fix $\alpha=0.3$ and search different configurations for the $\beta$. Then we fix $\beta$ to the optimal value in the Twitter2015 dev set and search for $\alpha$.
Parameter Analysis of $\beta$.
As shown in Fig. 8, we find that: 1) different models are sensitive to the values of $\beta$. 2) $\beta$ values within the range of $[0.4,0.5]$ demonstrate strong performance among various models. 3) Excessive values of $\beta$ result in performance degradation. We conjecture that this decline may be caused by introducing excessive noise within the context. For a fair comparison, we select the optimal value of $\beta$ on the dev set for evaluation.
Parameter Analysis of $\alpha$.
Fig. 9 shows that: 1) models exhibit insensitivity to the value of $\alpha$. 2) $\alpha$ values within the range of $[0.3,0.4]$ exhibit strong performance across diverse models. 3) There is no significant decrease in trends with increasing $\alpha$. Thus, we set $\alpha=0.3$ as our default setting.
B.2 Scalability of WisdoM
Our plug-in method is data- and model-agnostic, therefore, it is expected to be highly scalable. Here we scale our WisdoM[Uncaptioned image] up to different model sizes and data volumes.
Performance on Different Model Sizes.
We experiment with scaling the model size to see if there are ramifications when operating at a larger scale. Fig. 10 (a) reveals that the performance increases as the LVLM size increases. In addition, we found that as the size of the model increased, the performance gains became more pronounced.
Performance on different Data Volumes.
We conduct experiments on different ratios of training data to verify the robustness of WisdoM. As shown in Fig. 10 (b), we find that even when only 25% of the training data was used, our method WisdoM resulted in a 0.29% improvement.